# SmallToLarge (S2L): Scalable Data Selection for Fine-tuning Large Language Models by Summarizing Training Trajectories of Small Models

### v1

@@ Abstract
Despite the effectiveness of data selection for large language models (LLMs) during pretraining and instruction fine-tuning phases, improving data efficiency in supervised fine-tuning (SFT) for specialized domains poses significant challenges due to the complexity of fine-tuning data. To bridge this gap, we introduce an effective and scalable data selection method for SFT, SmallToLarge (S2L), which leverages training trajectories from small models to guide the data selection for larger models. We demonstrate through extensive experiments that S2L significantly improves data efficiency in SFT for mathematical problem-solving, reducing the training data to just 11% of the original MathInstruct dataset (Yue et al., 2023) to match full dataset performance while outperforming state-of-the-art data selection algorithms by an average of $4.7\%$ across 6 in- and out-domain evaluation datasets. Remarkably, selecting only 50K data for SFT, S2L achieves a 32.7% accuracy on the challenging MATH (Hendrycks et al., 2021) benchmark, improving Phi-2 (Li et al., 2023b) by 16.6%. In clinical text summarization on the MIMIC-III dataset (Johnson et al., 2016), S2L again outperforms training on the full dataset using only 50% of the data. Notably, S2L can perform data selection using a reference model 40×40$\times$40 × smaller than the target model, proportionally reducing the cost of data selection.
Machine Learning, ICML
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized artificial intelligence by demonstrating an unprecedented ability to understand and generate human language (Brown et al., 2020). Applications requiring a balanced performance across the full spectrum of natural text, such as open-ended dialogue (Thoppilan et al., 2022; Touvron et al., 2023) or instruction following (Ouyang et al., 2022; Wei et al., 2022a), favor generalist models. However, to maximize performance in a specialized domain, such as mathematics (Azerbayev et al., 2023; Luo et al., 2023a; Yu et al., 2023; Liu et al., 2023; Yue et al., 2023), code (Roziere et al., 2023; Luo et al., 2023b), medicine (Singhal et al., 2023a, b; Cheng et al., 2023), or finance (Wu et al., 2023a; Cheng et al., 2023), models fine-tuned on domain data offer superior capabilities over generalist models (Jang et al., 2023).
@@
(c) Hidden states of MathInstruct on fully fine-tuned Pythia-410M
(d) Increase in training time as the size of the model scales up
@@
Figure 1: Existing data selection methods often rely on the feature representation from a reference model, which makes their effectiveness very sensitive to how well they have been trained on the target dataset (Marion et al., 2023). For supervised fine-tuning (SFT), while the pretrained model can provide good feature representations for natural language and separate topics in different colors (Figure 0(a)), it cannot provide a good feature representation if the fine-tuning data has a distribution shift from the pretraining (Figure 0(b)), which is usually the case when the data is from specialized domains. Even after fine-tuning, the improvement in the feature representation is limited (LABEL:{fig:finetuned-math}). Meanwhile, the cost of training a reference model that can be used to get feature representation for data selection increases significantly as the model size increases, making the deployment of data selection for large language models expensive.
@@
Recent research indicates that LLMs benefit more from training for additional epochs on carefully curated data rather than on larger, uncurated ones during pretraining (Tirumala et al., 2023) and instruction fine-tuning (Zhou et al., 2023a), making data selection one of the most promising means to unlock the next level of LLMs’ language capability. Yet, maximizing the data efficiency in supervised fine-tuning (SFT) for specialized domains remains a challenge. Firstly, heuristic approaches that are effective in the instruction fine-tuning stage, like manual curation (Zhou et al., 2023a) or using advanced models such as GPT-4 for dataset evaluation (Chen et al., 2024), are less reliable due to the need for specialized knowledge and become costly with large volumes of uncurated fine-tuning data. Beyond these heuristic methods, other approaches rely on generating representations for each training example using a reference model, often utilizing metrics like perplexity (Marion et al., 2023), confidence (Swayamdipta et al., 2020; Varshney et al., 2022), or hidden states (Abbas et al., 2023; Tirumala et al., 2023; Yang et al., 2023c; Bhatt et al., 2024) as features. However, these techniques also fall short in SFT for specialized domains for two reasons: (1) the significant shift in data characteristics during SFT can render these traditional feature representations less informative (Figure 1), and (2) the computation and memory demands associated with generating representations for each training become prohibitive as these specialized domains often require larger models, some with up to 540 billion parameters (Chowdhery et al., 2023; Singhal et al., 2023a), leading to substantial scalability challenges (Figure 0(d)).
To tackle the challenges of data selection in SFT for specialized domains, we present SmallToLarge (S2L), a more effective and scalable data selection method. S2L operates by first gathering training trajectories for each training example using a smaller model. These trajectories are then clustered, and a balanced selection is made from these clusters. This method is grounded in the findings of recent research (Xia et al., 2023), which reveals consistent training dynamics across models of various sizes. Such consistency validates the use of smaller models as effective proxies for data selection in larger models. Additionally, S2L’s balanced sampling from clusters ensures data with all learning patterns, some of which might be missed in uniform or weighted sampling, are adequately represented.
@@
To validate S2L’s effectiveness, we applied it to the challenging tasks of SFT for mathematical problem-solving and clinical text summarization. Our experiments on MathInstruct (Yue et al., 2023) have shown that S2L can significantly reduce the required training data size to just 11% of the original dataset while still matching the performance levels of the full dataset, outperforming current state-of-the-art one-shot and online data selection algorithms by an average of 4.7% across 6 in- and out-domain evaluation datasets. Remarkably, on the MATH benchmark (Hendrycks et al., 2021), S2L attained a 32.7% accuracy with just 50K data points, improving the best open-sourced model under 3 billion parameters, Phi-2, by 16.6%. For clinical text summarization tasks on the MIMIC-III (Johnson et al., 2016) dataset, S2L outperforms training on the complete dataset, using only half the data. Unlike the existing methods that require training and getting features from large models, S2L achieves superior data efficiency using a model with as few as 70 million parameters, which is 40x smaller than the largest target model we train with 2.8 billion parameters.
@@ 2 Related Work
Foundations of Data Selection. Data selection has been well studied for small models and classification tasks. There are one-shot algorithms that select data based on rankings of the proposed training statistics, for example, the L2-norms of error and gradient vectors (EL2N and GraNd) (Paul et al., 2021), confidence and its variability across epochs (Swayamdipta et al., 2020), and the number of times each example is learned but then forgot at the subsequent training step (Toneva et al., 2019). Besides these heuristic indicators, there are embedding-based pruning algorithms (Sorscher et al., 2022) and online selection algorithms with theoretical performance guarantees for efficiency (Mirzasoleiman et al., 2020; Killamsetty et al., 2021a, b; Pooladzandi et al., 2022; Yang et al., 2023b) and robustness (Yang et al., 2022, 2023a; Deng et al., 2023). Coleman et al. proposed to use the intermediate feature representation of a small proxy model to select data for image classification. Most recently, data selection has shown great potential in more substantial training speedup when implemented on near-storage hardware (Prakriya et al., 2023), and data selection beyond supervised learning of image data, e.g., for self-supervised learning (Joshi & Mirzasoleiman, 2023) and multimodal learning (Abbas et al., 2023; Mahmoud et al., 2023), also emerged.
Data Efficient Training of Large Language Models. For the pre-training of LLMs, Marion et al. studied data quality indicators including Perplexity, Error L2-Norm (EL2N) (Paul et al., 2021), and memorization ranking (Biderman et al., 2023a), and found training on examples with middle Perplexity rankings outperforms training on examples selected based on the other two metrics, and sometimes even outperforms training on the entire dataset. Tirumala et al. uses pre-trained model embeddings to select data for LLM pre-training. The proposed algorithm, D4, first applies an embedding-based data de-duplication algorithm (Abbas et al., 2023) and then discards data points that are the closest to the K-Means cluster centroids in the embedding space (Sorscher et al., 2022) to ensure diversity.
@@
On fine-tuning LLMs, existing work on data efficiency primarily focused on manually curating high-quality instructions (Zhou et al., 2023a), or using strong closed-source models (e.g., GPT-4 (Achiam et al., 2023) or ChatGPT) to rate the quality of each training example (Eldan & Li, 2023; Li et al., 2023a; Chen et al., 2024). Bhatt et al. implemented an experimental design framework to evaluate the existing data selection methods for instruction fine-tuning of LLMs and found selecting facility locations based on hidden representations (i.e., embeddings) is the most effective. As the only data selection algorithm for specialized domains, SCIP (Yang et al., 2023c) focuses on pruning low-quality code data for training code LLMs. Since it relies on breaking the code syntax to understand the characteristics of low-quality code in the embedding (i.e, hidden states) space, adapting SCIP to domains other than Python code data is non-trivial.
@@
Adapting Large Language Models for Specialized Domains. The rapid development of large language models (LLMs) gives rise to new state-of-the-art models in specialized domains. For mathematical reasoning, Galactica (Taylor et al., 2022), MINERVA (Lewkowycz et al., 2022) and Llemma (Azerbayev et al., 2023) continue to train an LLM on large-scale math-related web data to improve a model’s general scientific reasoning; WizardMath (Luo et al., 2023a) and TinyGSM (Liu et al., 2023) fine-tune LLMs using supervised data. Similarly for medical LLMs, Cheng et al. continued training pre-trained LLMs on medical text, and (Singhal et al., 2023a, b) fine-tuned PaLM with instruction prompt tuning on medical domain data.
@@ 3 Problem Formulation
$\min_{\bm{\theta}}\mathcal{L}(\bm{\theta},D_{\text{train}})=-\frac{1}{n}\sum_{(\mathbf{x},\mathbf{y})_{i}\in D_{\text{train}}}\big{[}\log p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})\big{]}.$ (2)
Data Selection Objective. In a general setting for data selection, we consider a target language model $p_{\bm{\theta}}$ with parameters $\bm{\theta}$. Given a fixed data budget $B$, which constrains the number of data points that can be used for training, our objective is to select a subset $S\subseteq D_{\text{train}}$ that optimizes the performance of the target model on the full training set $D_{\text{train}}$. This objective can be formally expressed as:
$\begin{split}\min_{S\subseteq D_{\text{train}},|S|\leq B}&|S|\ \ \text{s.t.}\quad\min\mathcal{L}(\bm{\theta}_{s},D_{\text{train}})\end{split}$ (3)
where $\bm{\theta}_{s}=\arg\min_{\bm{\theta}}\mathcal{L}(\bm{\theta},S)$ is the model trained on the subset $S$. In practice, the subset $S$ is selected based a reference model $r_{\bm{\phi}}$ parameterized with $\bm{\phi}$, which generate representations for each data point $(\mathbf{x},\mathbf{y})_{i}\in D_{\text{train}}$ as $r_{\bm{\phi}}((\mathbf{x},\mathbf{y})_{i})$. These representations are utilized by a data selection algorithm to produce $S$.
In existing data selection algorithms, $\bm{\phi}$ is commonly either weights of the pre-trained target model or a target model that has been fully trained on the dataset $D_{\text{train}}$ with parameters $\hat{\bm{\theta}}=\arg\min_{\bm{\theta}}\mathcal{L}(\bm{\theta},D_{\text{train}})$. However, as evidenced by Figure 1, representations generated by either the pretrained or the fine-tuned model may not always be good enough for data in specialized domains that involve a data distribution shift towards unique notations and terminologies.
@@ 4 Methodology
Our proposed method, SmallToLarge (S2L), focuses on optimizing the data selection objective by leveraging training trajectories of training examples on a small model.
Training Trajectory.
Let $\bm{\phi}^{(t)}$ be the parameter of a small LM during the training on $D_{\text{train}}$ at time $t,t\in\mathcal{T}=\{1,...,T\}$. The representation of each data point used in S2L is a sequence of losses recorded during training the reference model $(r_{\bm{\phi}^{(t)}})_{t=1}^{T}$, where
$r_{\bm{\phi}^{(t)}}(\mathbf{x},\mathbf{y})=-\log p_{\bm{\phi}^{(t)}}(\mathbf{y}|\mathbf{x}),$ (4)
and $T$ is the length of the training trajectory. Note that $\bm{\phi}^{(t)}$ is trained for a fixed number of iterations from $\bm{\phi}^{(t-1)}$, thereby saved sequentially during the training of $\bm{\phi}$.
Cluster-based Data Selection.
Once the loss trajectories are recorded, we apply a clustering algorithm to group examples based on the similarity of their loss trajectories. This results in a set of clusters $\{C_{1},C_{2},\ldots,C_{k}\}$, where each cluster $C_{i}$ contains examples with similar loss behavior over the training epochs:
$\begin{split}C_{i}=\{(\mathbf{x},\mathbf{y})|i=\arg\min_{j}d(\mathbf{L}(\mathbf{x},\mathbf{y}),\mathbf{L}{C_{j}}&)\},\end{split}$ (5)
where $\mathbf{L}{(\mathbf{x},\mathbf{y})}=(r_{\bm{\phi}^{(1)}}(\mathbf{x},\mathbf{y}),\ldots,r_{\bm{\phi}^{(T)}}(\mathbf{x},\mathbf{y}))$ denotes the loss trajectory of the example $(\mathbf{x},\mathbf{y})$, $\mathbf{L}_{C_{i}}$ is the centroid of the loss trajectories in cluster $C_{i}$, and $d(\cdot,\cdot)$ is a distance metric, such as Euclidean distance, used for clustering. As shown in Figure 2, clustering algorithms can effectively find groups of examples with similar training dynamics. With the clusters formed, the data selection strategy prioritizes selecting examples from smaller clusters while equally selecting examples from larger clusters, as detailed in Algorithm 1.
@@
Figure 2: Training trajectories ($r_{\bm{\phi}^{(t)}}(\mathbf{x},\mathbf{y})$) of 5 random examples in the same cluster vs. in different random clusters with k-means clustering. Examples in the same clusters have very similar training trajectories (Figure 1(a)) while the training trajectories of examples in different clusters are very different (Figure 1(b)).
Algorithm 1 Data Selection Based on Training Trajectories
@@
0:  Subset $S$ optimizing Equation 3.
@@
2:  Cluster examples in $D_{\text{train}}$ based on their training trajectories to form $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{K}\}$, sorted by size.
@@
8:        $S\leftarrow S\bigcup S_{k}$, where $S_{k}\subset C_{k},|S_{k}|=R_{k}$
@@
Intuition Behind Loss Trajectory Clustering.
Loss trajectories encapsulate information about how the model learns from specific data points. Consider two training examples $(\mathbf{x},\mathbf{y})_{i}$ and $(\mathbf{x},\mathbf{y})_{j}$ from the dataset $D_{\text{train}}$. If when the model effectively learns from example $(\mathbf{x},\mathbf{y})_{i}$, reflected in a decreasing loss trajectory $r_{\bm{\phi}^{(t)}}(\mathbf{x},\mathbf{y})_{i}$ over iterations, a correlated decrease in $r_{\bm{\phi}^{(t)}}(\mathbf{x},\mathbf{y})_{j}$ is observed, it is likely the two examples encode the same knowledge or skills at the same difficulty level. Therefore, by identifying clusters of examples with similar training trajectories and sampling examples from different clusters to cover all topics and difficulties, one can effectively train the model with fewer data points without significantly compromising the training quality. In Section 5.2.3 (Figure 5), we will show quantitatively such similarity between examples in the same cluster using the MathInstruct (Yue et al., 2023) dataset.
Small-to-Large Data Selection.
Training a reference model that is of the same size as the target model and getting the feature representation for each example in $D_{\text{train}}$ can be costly when the size of the target model is large. Inspired by a recent finding that the training dynamic of most examples is consistent across models of various sizes (Xia et al., 2023), in Figure 3, we empirically show that we can find groups of examples with similar training dynamics on larger models by clustering the training trajectories of $D_{\text{train}}$ on a small model, even when the trends are different.

### v2

# SmallToLarge (S2L): Scalable Data Selection for Fine-tuning Large Language Models by Summarizing Training Loss Trajectories of Small Models

@@ Abstract
Despite the effectiveness of data selection for pretraining and instruction fine-tuning large language models (LLMs), improving data efficiency in supervised fine-tuning (SFT) for specialized domains poses significant challenges due to the complexity of fine-tuning data. To bridge this gap, we introduce an effective and scalable data selection method for SFT, SmallToLarge (S2L), which trains a small model, clusters loss trajectories of the examples, and samples from these clusters to guide data selection for larger models. We prove that during fine-tuning, samples within the same loss trajectory cluster exhibit similar gradients. Then, we show that S2L subsets have a bounded gradient error w.r.t. the full data, hence guarantee convergence to the neighborhood of the optimal solution. We demonstrate through extensive experiments that S2L significantly improves data efficiency in SFT for mathematical problem-solving, reducing the training data requirement to just 11% of the original MathInstruct dataset [64] to match full dataset performance while outperforming state-of-the-art data selection algorithms by an average of $4.7\%$ across 6 in- and out-domain evaluation datasets. Remarkably, selecting only 50K data for SFT, S2L achieves a 32.7% accuracy on the challenging MATH [19] benchmark, improving Phi-2 [28] by 16.6%. In clinical text summarization on the MIMIC-III dataset [21], S2L again outperforms training on the full dataset using only 50% of the data. Notably, S2L can perform scalable data selection using a reference model 100×100$\times$100 × smaller than the target model, proportionally reducing the computational cost.  [^1] Code is available at https://github.com/BigML-CS-UCLA/S2L.
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized artificial intelligence by demonstrating an unprecedented ability to understand and generate human language [7]. Among all the contributing factors, the quality and selection of data is becoming increasingly recognized for its importance in training LLMs effectively. Recent research indicates that LLMs benefit more from training for additional epochs on carefully curated data rather than on larger, uncurated ones during pretraining [48] and instruction fine-tuning [67], making data selection one of the most promising means to unlock the next level of LLMs’ language capability.
@@
(c) Increase in training time as the size of the model scales up
@@
Figure 1: Existing data selection methods depend heavily on the feature representations from a reference model, which makes their effectiveness vulnerable to the quality of training on the target domain [34]. For supervised fine-tuning (SFT), while pretrained models can effectively separate topics (shown in different colors) in natural language (Figure 1(a)), they struggle with fine-tuning data that deviates from the pretraining distribution (Figure 1(b)). Additionally, the cost of training a reference model escalates with model size (Figure 1(c)), making existing data selection methods for large models prohibitively expensive.
However, while generalist models obtained through pre-training or instruction fine-tuning excel in general language tasks, they may not deliver optimal outcomes in specialized domain, such as mathematics [3, 31, 63, 30, 64], code [42, 32], medicine [43, 44, 9], or finance [57, 9]. These domains are not only critical for real-world applications but also hold substantial economic and societal impacts.
@@
To maximize performance in specialized domains, models fine-tuned on domain data offer superior capabilities over generalist models [20]. Yet, maximizing the data efficiency in supervised fine-tuning (SFT) for specialized domains remains a challenging and under-explored problem. Firstly, heuristic approaches that are effective in the instruction fine-tuning stage, like manual curation [67] or using advanced models such as GPT-4 for dataset evaluation [8], are less reliable due to the need for specialized knowledge and become costly with large volumes of uncurated fine-tuning data. Beyond these heuristic methods, other approaches rely on generating representations for each training example using a reference model, often utilizing metrics like perplexity [34], confidence [46, 53], or hidden states [1, 48, 61, 4] as features. However, these techniques also fall short in SFT for specialized domains for two reasons: (1) the significant shift between pretraining and SFT data can render these metrics less informative (Figure 1(b)), and (2) the computation and memory demands associated with generating representations for each training example become prohibitive, as these specialized domains often require larger models, some with up to 540 billion parameters [10, 43], leading to substantial scalability challenges (Figure 1(c)).
To tackle the challenges of data efficiency in SFT for specialized domains, we present SmallToLarge (S2L), an effective and scalable data selection algorithm. S2L operates by first gathering training loss trajectories for each training example using a small model. These trajectories are then clustered, and similar number of examples are selected from these clusters uniformly at random. This process is grounded in our theoretical findings that examples within the same cluster exhibit similar gradients during training, thereby affecting the model similarly. Consequently, subsets sampled from these clusters have a bounded gradient error w.r.t. the full data, allowing for training a comparable model with only a subset of data. Furthermore, we provide a convergence rate analysis for training on these subsets, establishing a robust theoretical foundation for S2L’s effectiveness and efficiency.
@@
To validate S2L’s effectiveness, we applied it to the challenging tasks of SFT for (1) mathematical problem-solving and (2) clinical text summarization. Our experiments on MathInstruct [64] shows that S2L can significantly reduce the required training data size to just 11% of the original dataset size while still matching the performance levels of the full dataset, outperforming current state-of-the-art one-shot and online data selection algorithms by an average of 4.7% across 6 in- and out-domain evaluation datasets. Remarkably, on the MATH benchmark [19], S2L attained a 32.7% accuracy with just 50K data points, improving the best open-sourced model under 3 billion parameters, Phi-2, by 16.6%. For clinical text summarization tasks on the MIMIC-III [21] dataset, S2L outperforms training on the full dataset, using only half of the data. Unlike existing methods that require training and getting features from large models, S2L achieves superior data efficiency using a model with as few as 70 million parameters, which is 100×100$\times$100 × smaller than the largest target model we train with 7 billion parameters.
@@ 2 Related Work
Foundations of Data Selection. Data selection has been well studied for small models and classification tasks. There are one-shot algorithms that select data based on rankings of the proposed training statistics, for example, the L2-norms of error and gradient vectors (EL2N and GraNd) [39], confidence and its variability across epochs [46], and the number of times each example is learned but then forgot at the subsequent training step [49]. Besides these heuristic indicators, there are embedding-based pruning algorithms [45] and online selection algorithms with theoretical performance guarantees for efficiency [35, 23, 24, 40, 60] and robustness [59, 62, 16]. Coleman et al. proposed to use the intermediate feature representation of a small proxy model to select data for image classification. Most recently, data selection has shown great potential in more substantial training speedup when implemented on near-storage hardware [41], and data selection beyond supervised learning of image data, e.g., for self-supervised learning [22] and multimodal learning [1, 33], also emerged.
@@
Data Efficient Training of Large Language Models. For the pre-training of LLMs, Marion et al. studied data quality indicators including Perplexity, Error L2-Norm (EL2N) [39], and memorization ranking [5], and found training on examples with middle Perplexity rankings outperforms training on examples selected based on the other two metrics, and sometimes even outperforms training on the entire dataset. Tirumala et al. uses pre-trained model embeddings to select data for LLM pre-training. The proposed algorithm, D4, first applies an embedding-based data de-duplication algorithm [1] and then discards data points that are the closest to the K-Means cluster centroids in the embedding space [45] to ensure diversity. On fine-tuning LLMs, existing work on data efficiency primarily focused on manually curating high-quality instructions [67], or using strong closed-source models (e.g., GPT-4 [2] or ChatGPT) to rate the quality of each training example [18, 27, 8]. Bhatt et al. implemented an experimental design framework to evaluate the existing data selection methods for instruction fine-tuning of LLMs and found selecting facility locations based on hidden representations (i.e., embeddings) is the most effective. As the only data selection algorithm for specialized domains, SCIP [61] focuses on pruning low-quality code data for training code LLMs. Since it relies on breaking the code syntax to understand the characteristics of low-quality code in the embedding (i.e, hidden states) space, adapting SCIP to domains other than Python code data is non-trivial.
@@
Adapting Large Language Models for Specialized Domains. The rapid development of large language models (LLMs) gives rise to new state-of-the-art models in specialized domains. For mathematical reasoning, Galactica [47], MINERVA [26] and Llemma [3] continue to train an LLM on large-scale math-related web data to improve a model’s general scientific reasoning; WizardMath [31] and TinyGSM [30] fine-tune LLMs using supervised data. Similarly for medical LLMs, Cheng et al. continued training pre-trained LLMs on medical text, and [43, 44] fine-tuned PaLM with instruction prompt tuning on medical domain data.
@@ 3 Problem Formulation
$\min_{\bm{\theta}}\mathcal{L}(\bm{\theta},D_{\text{train}})=-\frac{1}{n}\sum_{(\mathbf{x},\mathbf{y})_{i}\in D_{\text{train}}}\big{[}\log p_{\bm{\theta}}(\mathbf{y}_{i}|\mathbf{x}_{i})\big{]}.$ (2)
Data Selection Objective. In a general setting for data selection, we consider a target language model $p_{\bm{\theta}}$ with parameters $\bm{\theta}$. Given a fixed data budget $B$, which constrains the number of data points that can be used for training, our objective is to select a subset $S\subseteq D_{\text{train}}$ to train the target model, such that it obtains a superior generalization performance. In practice, the subset $S$ is selected based on a reference model $r_{\bm{\phi}}$ parameterized by $\bm{\phi}$, which generates representations, confidence scores, or other metrics for each data point $(\mathbf{x},\mathbf{y})_{i}\in D_{\text{train}}$, denoted by $r_{\bm{\phi}}((\mathbf{x},\mathbf{y})_{i})$, which will be utilized by a data selection algorithm to produce $S$.
In existing data selection algorithms, $\bm{\phi}$ is commonly either weights of the pre-trained target model or a target model that has been fully trained on the dataset $D_{\text{train}}$. However, as evidenced by Figure 1, representations generated by the pretrained model may not always be good enough for data selection in specialized domains, and fine-tuning the target model significantly increases the computational cost of data selection.
@@ 4 Methodology
Training a large target model to obtain feature representations for each example in $D_{\text{train}}$ can be computationally intensive. However, a recent finding demonstrates that the training dynamics of most examples are consistent across differently sized models of the same family, and this phenomena even generalizes across different model families [58]. Our proposed method, SmallToLarge (S2L), leverages loss trajectories of training examples collected during fine-tuning a small reference model on the full or a subset of training data.
Loss Trajectory.
Let $\bm{\phi}^{(t)}$ be the parameters of a small LM during training on $D_{\text{train}}$ at times $t_{q},q\in\{1,...,T\}$. S2L records the loss trajectory for each data point $i$ at times $t_{q}$ during training the reference model $[\mathcal{L}_{i}^{\text{proxy}}(\bm{\phi}^{(t_{1})}),\ldots,\mathcal{L}_{i}^{\text{proxy}}(\bm{\phi}^{(t_{T})})]$ where
$\begin{split}\mathcal{L}_{i}^{\text{proxy}}(\bm{\phi}^{(t)})=\mathcal{L}^{\text{proxy}}(\bm{\phi}^{(t)},(\mathbf{x}_{i},\mathbf{y}_{i}))=-\log p_{\bm{\phi}^{(t)}}(\mathbf{y}_{i}|\mathbf{x}_{i}),\end{split}$ (3)
and $T$ is the length of the loss trajectory. Note that $\bm{\phi}^{(t)}$ is trained for a fixed number of iterations from $\bm{\phi}^{(t-1)}$.
Assume the parameter vector $\bm{\theta}^{(t)}$ represents the parameters of the target model at the time $t$. Define $\mathbf{L}_{i}^{\text{proxy}}=[\mathcal{L}_{i}^{\text{proxy}}(\bm{\phi}^{(t_{1})}),\ldots,\mathcal{L}_{i}^{\text{proxy}}(\bm{\phi}^{(t_{T})})]$ and $\mathbf{L}_{i}^{\text{target}}=[\mathcal{L}_{i}^{\text{target}}(\bm{\theta}^{(t_{1})}),\ldots,\mathcal{L}_{i}^{\text{target}}(\bm{\theta}^{(t_{T})})]$ as the training loss trajectory of the example $i$ on the small proxy model and the large target model, respectively. Let $\bm{H}_{i}\in\mathbb{R}^{d\times d}$ be the Hessian matrix for each example $i$ and assume that the loss function for each example during fine-tuning can be modeled by a second-order Taylor approximation with bounded curvature ($c\leq\|\bm{H}_{i}\|\leq C$), a reasonable assumption in fine-tuning settings. The following lemma shows that examples with similar loss trajectories on the proxy model have similar gradients throughout the training of the target model.
Theorem 4.1.
If examples $i$ and $j$ have similar loss trajectories on the proxy model, i.e., $\|\mathbf{L}_{i}^{\text{proxy}}-\mathbf{L}_{j}^{\text{proxy}}\|\leq\epsilon$, and their loss trajectories on the proxy and target model is similar, i.e., $\|\mathbf{L}_{p}^{\text{proxy}}-\mathbf{L}_{p}^{\text{target}}\|\leq\delta$ for $p\in\{i,j\}$, then $i$ and $j$ have similar gradients throughout training the target model:
$\|\nabla\mathcal{L}_{i}^{\text{target}}(\bm{\theta})-\nabla\mathcal{L}_{j}^{\text{target}}(\bm{\theta})\|\leq\frac{2\epsilon^{\prime}+2CD^{2}}{d}=\Delta.$ (4)
where $\epsilon^{\prime}=\epsilon+2\delta$ and $\|\bm{\theta}\|\leq D$ for all $t$.
The proof of Theorem 4.1 can be found in Section A.1. Theorem 4.1 shows that examples with similar loss trajectories have similar gradients during the training, thereby influencing the model in a similar manner.
@@
Figure 2: Examples in the same clusters have very similar loss trajectories (Figure 2(a)) while the loss trajectories of examples in different clusters are very different (Figure 2(b)).
Data selection from Loss Trajectory Clusters.
Once the loss trajectories are recorded on the proxy model, we apply a clustering algorithm to group examples based on the similarity of their loss trajectories. This results in a set of clusters $\{C_{1},C_{2},\ldots,C_{K}\}$, where each cluster $C_{i}$ contains examples with similar loss and gradient trajectory throughout the training:
$\begin{split}C_{i}=\{(\mathbf{x},\mathbf{y})_{j}\in D_{\text{train}}|i=\arg\min_{j\in[K]}d(\mathbf{L}_{j},\mathbf{L}_{\bar{C_{j}}}&)\},\end{split}$ (5)
where $\mathbf{L}_{\bar{C_{i}}}$ is the centroid of the loss trajectories in cluster $C_{i}$, and $d(\cdot,\cdot)$ is a distance metric, such as Euclidean distance, used for clustering. For datasets that contain different sources of data, we cluster each source separately.
Algorithm 1 Data Selection Based on Training Trajectories (S2L)
@@
0:  Subset $S\subseteq D_{\text{train}},|S|\leq B$.
@@
2:  Train a small proxy model and cluster examples in (each data source of) $D_{\text{train}}$ based on their loss trajectories and sort them by size to get $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{K}\}$.
@@
8:        $S\leftarrow\{S\bigcup S_{k}$}, where $S_{k}\subset C_{k}$ is selected uniformly at random from $C_{k}$ and $|S_{k}|=R_{k}$
@@
As shown in Figure 3, clustering algorithms can effectively find groups of examples with similar training dynamics. In Figure 3, we empirically show that we can identify groups of examples with similar training dynamics on a larger model by clustering the training trajectories of $D_{\text{train}}$ on a smaller proxy model. With the clusters formed, the data selection strategy selects equal number of examples at random from all clusters, as detailed in Algorithm 1. In doing so, it effectively prioritizes selecting examples from smaller clusters. This is particularly important for datasets containing multiple imbalanced sources. In this setting, training and test distributions often differ, and balanced selection from clusters ensures superior test performance on all groups in the test data.
The following theorem shows that, under the assumptions of Theorem 4.1, training with Incremental Gradient (IG) methods on the subset selected by S2L converges to a close neighborhood of the optimal solution found by training the target model on the full dataset. IG methods such as Stochastic Gradient Descent (SGD) update parameters iteratively based on the gradient of the loss of individual examples, multiplied by stepsize $\alpha$. Formally,
$\bm{\theta}^{t+1}=\bm{\theta}^{t}-\alpha\nabla\mathcal{L}_{i}^{\text{target}}(\bm{\theta}^{t}).$ (6)
Corollary 4.2.
Under the assumptions of Theorem 4.1, applying IG with stepsize $\alpha$ to subsets found by S2L, converges to the neighborhood of the optimal solution, as follows:
$\|\bm{\theta}^{t+1}-\bm{\theta}^{*}\|^{2}\leq(1-\alpha c)^{t+1}\|\bm{\theta}^{t}-\bm{\theta}^{*}\|^{2}+2\xi R/c^{2}+\alpha B^{2}(r_{\min}/k)^{2}\bm{g}_{\max}^{2}$ (7)
where $c\leq\|\bm{H}\|$, $B=k\cdot K$ is the total size of the subset, $\bm{g}_{\max}$ is the largest gradient norm of individual examples during training, $r_{\min}=\min_{j}|C_{j}|,r_{\max}=\max_{j}|C_{j}|$, $R=\min\{d_{0},B\bm{g}_{\max}+\xi/c\}$ and $d_{0}=\|\bm{\theta}^{0}-\bm{\theta}^{*}\|$ is the initial distance to the optimal solution $\bm{\theta}^{*}$, and $\xi$ is given by:
$\xi=K[r_{\min}\Delta+(r_{\max}-r_{\min})\bm{g}_{\max}].$ (8)
The proof can be found in Section A.2.